The horizontal coordinates of a Frisbee in a strong wind are given by ? = −12 + 4? 3 and ? = 10? − 3? 2 where x and y are in meters, and t is in seconds. A. Find the average velocity of the Frisbee between ? = 0? and ? = 2?. (6 marks) B. Find the velocity of the Frisbee at ? = 0?. (2 marks) C. Find the acceleration of the Frisbee at ? = 1?. (2 marks)Question 4: (10 Marks) The horizontal coordinates of a Frisbee in a strong wind are given by x = -12 + 4t3 and y = 10t – 3t2
Answer #1

x=-12 + 4 +3 y = lot-342 at os at ~ Ls a = -12 + 4x 23 2om y = tox2-3x23 8. = (120) (20,8) Displace |(20+1242 +82 32 1981 Var